OpenStudy (anonymous):

Factor: x2 - 9x + 12

OpenStudy (anonymous):

Well, the outer terms have to equal 12. The inner terms have to add to equal -9. What options do you have?

OpenStudy (anonymous):

A. (x + 1)(x - 12) B. (x + 2)(x + 6) C. (x - 2)(x - 6) D. prime

OpenStudy (anonymous):

It's prime. None of the other options produce -9x when the inside and outside terms are added together.... Does that make sense to you?

OpenStudy (anonymous):

yes, thank you (:

OpenStudy (unklerhaukus):

\[(x + 1)(x - 12)=x^2-11x-12\neq x^2 - 9x + 12\] \[(x + 2)(x + 6)=x^2+8x+12\neq x^2 - 9x + 12\] \[(x - 2)(x - 6)=x^2-8x+12\neq x^2 - 9x + 12\]

OpenStudy (anonymous):

\[x=\frac{-b^2\pm \sqrt{b^2-4ac}}{2a}\]a=1, b= -9 , c=12